### v1

# Test-time regression: a unifying framework for designing sequence models with associative memory

Abstract
Sequences provide a remarkably general way to represent and process information. This powerful abstraction has placed sequence modeling at the center of modern deep learning applications, inspiring numerous architectures from transformers to recurrent networks. While this fragmented development has yielded powerful models, it has left us without a unified framework to understand their fundamental similarities and explain their effectiveness. We present a unifying framework motivated by an empirical observation: effective sequence models must be able to perform associative recall. Our key insight is that memorizing input tokens through an associative memory is equivalent to performing regression at test-time. This regression-memory correspondence provides a framework for deriving sequence models that can perform associative recall, offering a systematic lens to understand seemingly ad-hoc architectural choices. We show numerous recent architectures — including linear attention models, their gated variants, state-space models, online learners, and softmax attention — emerge naturally as specific approaches to test-time regression. Each architecture corresponds to three design choices: the relative importance of each association, the regressor function class, and the optimization algorithm. This connection leads to new understanding: we provide theoretical justification for QKNorm in softmax attention, and we motivate higher-order generalizations of softmax attention. Beyond unification, our work unlocks decades of rich statistical tools that can guide future development of more powerful yet principled sequence models.
Section 1 Introduction
Sequences play a vital role in modern machine learning by providing a powerful abstraction: any computational task can be viewed as transforming one sequence into another (Sutskever et al., 2014). This sequential perspective has spread across diverse domains, including natural language processing (Sutskever et al., 2014, Devlin et al., 2019, Brown et al., 2020), computer vision (Dosovitskiy et al., 2021, Bertasius et al., 2021), time series analysis (Salinas et al., 2020, Gruver et al., 2023, Ansari et al., 2024), and computational biology (Jumper et al., 2021, Zhou and Troyanskaya, 2015, Nguyen et al., 2024), highlighting the importance of building generically applicable sequence layers (Vaswani et al., 2017).
This development has produced a diversity of architectures, each with its own unique characteristics and performance trade-offs. While these architectures have achieved considerable success, they have largely emerged through separate lines of investigation. This fragmented and often empirically-driven approach to model development limits our ability to systematically understand and improve design choices. Moreover, the idiosyncratic notations of each architecture obscures their underlying connections (Rush, 2024). Given the wide variety of sequence models, from transformers to convolutional networks to recurrent networks, a natural question to ask is whether there is an underlying principle that explains why some sequence models work better than others.
One empirical discovery that ties together disparate architectures is the strong correlation between their associative recall ability and their language modeling performance (Olsson et al., 2022, Arora et al., 2023). For example, Olsson et al. (2022) found that transformers learn induction heads that store consecutive tokens as key-value pairs to predict future tokens. Given the empirical importance of associative recall, how can we systematically design architectures that can perform associative recall? In this paper, we introduce a simple but principled framework that results in sequence models with associative memory, unifying broad classes of models.
Our crucial observation is that we can memorize key-values pairs by solving a regression problem (Kohonen, 1972). A sequence layer that memorizes earlier tokens for later retrieval is one that performs test-time regression over the input tokens; we name this “test-time regression”, a reference to the test-time training paradigm (Sun et al., 2024, 2020). By establishing the equivalence between associative memory and regression, we create a framework under which we can holistically understand ad-hoc architecture developments and motivate new advances. In particular, our framework allows us to repurpose well-understood regression tools, including weighted least-squares regression, online algorithms for regression, and nonparametric regression to explain existing architectures. Under our framework, a sequence layer is a mathematical consequence of choosing a regression objective, a regressor parameterization, and an optimization algorithm.
To illustrate the generality of our framework, we concisely derive linear attention (Katharopoulos et al., 2020), its feature-mapped variants (e.g. (Peng et al., 2020, Qin et al., 2021, Kasai et al., 2021, Zhang et al., 2023, Aksenov et al., 2024, Chen et al., 2024)), its gated variants (Sun et al., 2023, Orvieto et al., 2023, Katsch, 2024, De et al., 2024, Qin et al., 2024, Peng et al., 2024, Yang et al., 2024b, Beck et al., 2024), state-space models (Gu and Dao, 2024, Dao and Gu, 2024), fast-weights programmers (Schlag et al., 2021, Yang et al., 2024a, a), online learners (Liu et al., 2024, Sun et al., 2024, Yang et al., 2024a, Behrouz et al., 2024), and even softmax self-attention (Vaswani et al., 2017) as points within our design space. Past work by Schmidhuber (1992), Schlag et al. (2021), Clark et al. (2022), Liu et al. (2024), Yang et al. (2024a), Sun et al. (2024) proposed test-time training/fast-weight architectures that perform gradient descent at test time. Concurrent work by Behrouz et al. (2024) also points out the relationship between online learners and memorization. In contrast, we show that test-time regression is an even more fundamental principle that can also derive other classes of sequence layers, including the aforementioned gated linear attention models, state space models, softmax attention, and linear regression layers (von Oswald et al., 2023, Garnelo and Czarnecki, 2023). Our framework abstracts away practically important but low-level details in exchange for a clearer perspective on what these architectures have in common, enabling direct comparison. The regression-memory correspondence unlocks decades of advances in classical regression theory (Hastie et al., 2009, Green and Silverman, 1993, Wasserman, 2006), offering a systematic path toward sequence architectures that are both more powerful and theoretically sound.
Figure 1: Our test time regression framework unifies existing sequence models through the regression-memory correspondence. The equivalence between in-context regression and associative memory leads to a systematic approach for sequence model design through three key choices: the relative importance of each association, the regressor function class, and the optimization algorithm. The framework encompasses models ranging from classical linear regression with exact solvers (left) to first-order optimization approaches (middle), to nonlinear regressors (right). This unification reveals that seemingly different architectural innovations (shown in italics) are instances of well-studied concepts from regression theory but implemented as a neural network layer.
Section 1.1 Outline of our paper
In Section 2, we introduce our test-time regression framework and formalize the connection between associative memory and regression. We show how this perspective provides a systematic approach to sequence model design through three key choices: regression objective, function class, and optimization algorithm. Section 3 demonstrates the broad applicability of our framework by deriving state-of-the-art sequence architectures from first principles. In the process, we explain the effectiveness of QKNorm in softmax attention, and we derive higher-order generalizations of softmax attention, among other contributions. Section 4 examines how to construct effective key-value pairs for associative recall. Finally, Section 5 discusses the broader implications of viewing sequence models through the lens of regression, including future directions for new architectures.
Section 2 Test-time regression as a framework for model design
The goal of sequence modeling is to transform a sequence of input tokens $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ into output tokens $\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ (Sutskever et al., 2014). Although many sequence layers exist, past works have shown that associative recall ability is crucial to this task by enabling in-context learning (Olsson et al., 2022). Though many sequence layers exist, it is not a priori clear which ones are able to perform associative recall without empirical evaluations (Arora et al., 2023). A common design choice is to transform the input tokens into key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$ and queries $\mathbf{q}_{1},\ldots,\mathbf{q}_{T}$ before mixing them, following the success of transformers (Vaswani et al., 2017). While transforming inputs into query-key-value tokens has proven empirically effective, the field has lacked theoretical principles for designing sequence layers with associative memory capabilities. We address this gap through our test-time regression framework, which provides a principled recipe for deriving sequence architectures with mathematically grounded recall abilities.
Associative memory as regression.
From our everyday experience, we already have an intuition for what an associative memory should behave like; for example, hearing a friend’s name should trigger a mental impression of that friend. This pairing is called a “cue” and a “response”. To reflect terminology in deep learning (Vaswani et al., 2017), we will refer to cues as “keys” and responses as “values”. Given a set of associations $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$, an associative memory system is a system that performs associative recall: return $\mathbf{v}_{i}$ when given $\mathbf{k}_{i}$. Such a mapping between different sets of related objects is known as a hetero-associative memory in classic signal processing and neurocomputing literature (Hinton and Anderson, 1989).
With this intuition, we can now define a mathematical model for associative memory. Given key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})\in\mathbb{R}^{D_{k}}\times\mathbb{R}^{D_{v}}$ to be memorized, an associative memory system is a function ${m}:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{v}}$ such that ${m}(\mathbf{k}_{t})\approx\mathbf{v}_{i}$ for $t=1,\ldots,T$ (Kohonen, 1972, 1989). When ${m}$ is flexible enough to interpolate all values such that ${m}(\mathbf{k}_{t})=\mathbf{v}_{t}$ for all $t$, it is a perfect associative memory system. This mapping implements a form of fuzzy memory: querying with $\mathbf{q}$ that approximates or is a corrupted version of some key $\mathbf{k}_{i}$ will retrieve a value close to $\mathbf{v}_{i}$, making the memory robust to input noise and perturbations (Hinton and Anderson, 1989).
Our key observation is that finding a memory map that satisfies these constraints reduces to finding a vector-valued regressor ${m}$. We can derive an associative memory system by solving the weighted regression problem
$\operatorname*{minimize}\limits_{{m}\in\mathcal{M}}\ \ \frac{1}{2}\sum_{i=1}^{T}\gamma_{i}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2},$ (1)
where the importance of each association is controlled by adjusting the relative weights $\{\gamma_{i}\}_{i=1}^{T}$ Classic works on associative memory are restricted to linear memory maps, corresponding to linear regression (Kohonen, 1989, Chapter 6.6), but here we allow $\mathcal{M}$ to be any function class. We refer to this relationship between regression and associative memory as the “regression-memory correspondence”.
Having established regression as a way to implement associative memory, we design sequence layers by solving Equation 1 to construct memory maps ${m}_{A}$ that satisfy ${m}_{A}(\mathbf{k}_{i})\approx\mathbf{v}_{i}$ for associations $A=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i\in I}$. For non-causal tasks (e.g. vision transformers), we construct a single memory ${m}_{A}$ over all key-value pairs and compute outputs as $\mathbf{y}_{t}={m}_{A}(\mathbf{q}_{t})$. For causal tasks (e.g. language models), we construct separate memories ${m}_{t}$ for each prefix $A_{t}=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{t}$ and compute outputs as $\mathbf{y}_{t}={m}_{t}(\mathbf{q}_{t})$. The forward pass thus solves Equation 1 to obtain ${m}$ before applying it to queries. This naturally extends to multihead architectures (Vaswani et al., 2017, Shazeer, 2019, Ainslie et al., 2023), which correspond to parallel associative memories that may share keys, values, or queries.
We call such layers “test-time regression layers”, referencing the similar idea of test-time-training (Sun et al., 2020) but specialized to associative memory. Though solving Equation 1 can be seen as a self-supervised test-time training objective that regresses one view upon another (Sun et al., 2024), our associative memory perspective is more intuitive given its importance in in-context learning (Arora et al., 2023, Olsson et al., 2022).
A recipe for designing your own architecture.
Although Equation 1 succinctly summarizes our framework, it leaves many details unspecified. For example, over which class of regressors $\mathcal{M}$ should we consider? Additionally, which associations should we prioritize memorizing? Finally, which optimization procedure should we use to minimize the loss? The regression-memory correspondence thus allows for a vast design space. We partition this wide landscape into three simple design choices:
  1. 1.
    the relative importance of each association, specified through weights in Equation 1,
  2. 2.
    the function class $\mathcal{M}$ over which we search,
  3. 3.
    the minimization algorithm,
giving us a formulaic “recipe” for deriving sequence layers that can perform associative recall.
Section 3 Deriving existing architectures from regression
In this section, we consider causal sequence models that construct a new ${m}_{t}$ at each timestep using associations $A_{t}=\{(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{t},\mathbf{v}_{t})\}$. The case for non-causal sequence models easily generalizes, following our previous discussion.
To demonstrate how our framework unifies our understanding of sequence models, we succinctly derive a swath of architectures commonly studied today. We categorize existing sequence layers into four distinct classes: linear attention (possibly with feature maps), linear attention with gating (which includes state-space models like Mamba), online learners, and softmax self-attention.
Notation.
We first explain the mathematical notation we will use in our derivations. We use bold lowercase letters to refer to vectors (e.g. $\mathbf{x},\mathbf{k},\mathbf{v}$) and bold uppercase letters to refer to matrices (e.g. $\mathbf{M},\mathbf{K},\mathbf{V}$). Assume we have a sequence of $T$ inputs $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ which have already transformed into a sequence of keys and values $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$. We will use these keys and values as the dataset for our test-time regression problem. Let $\mathbf{K}_{t}=\begin{bmatrix}\mathbf{k}_{1}\ldots\mathbf{k}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{k}}}$ and $\mathbf{V}_{t}=\begin{bmatrix}\mathbf{v}_{1}\ldots\mathbf{v}_{t}\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{v}}}$ be matrices that contain $t$ keys and $t$ values respectively, for $t=1,\ldots,T$. Each key/value vector is a row of the key/value matrix. To minimize clutter, we will use the shorthand $\mathbf{K}$ and $\mathbf{V}$ instead of $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ when the timestep $t$ is clear from context.
Vignette 1: Linear attention is (suboptimal) linear least squares regression
We start by deriving an architecture from linear least squares, the most classic regression method. Linear least squares corresponds to the following choices in our recipe.
  1. 1.
    Choice of weights: assign equal weight to each association,
    ${m}_{\textsc{linear},\,t}=\operatorname*{\arg\!\min}_{{m}\in{\mathcal{M}_{\textsc{linear}}}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2}.$ (2)
  2. 2.
    Choice of function class: linear functions ${\mathcal{M}_{\textsc{linear}}}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\mathbf{k},\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}\}$.
  3. 3.
    Choice of minimization algorithm: analytical solution equivalent to one step of Newton’s method [^3] Since the objective function is quadratic, the analytical solution is equivalent to one iteration of Newton’s method initializing from the origin..
Since ${m}_{\textsc{linear},\,t}$ is a linear function, it is equivalent to a linear map $\mathbf{M}_{t}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}$. Minimizing the linear least squares objective results in the solution:
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\mathbf{k}_{i}\rVert_{2}^{2}=\begin{cases}\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1},\ \text{overdetermined, e.g. }t\geq{D_{k}},\\ \mathbf{V}{}^{\top}(\mathbf{K}\mathbf{K}{}^{\top})^{-1}\mathbf{K},\ \text{underdetermined, e.g. }t<{D_{k}},\end{cases}$ (3)
where we choose the min-norm solution in the underdetermined case. The overdetermined case corresponds to $\mathbf{K}$ having linearly-independent columns, which is not possible if $t<{D_{k}}$, while the underdetermined case corresponds to $\mathbf{K}$ having linearly-independent rows, which is not possible if $t>{D_{k}}$. We restrict our discussion to the case when $t\geq{D_{k}}$, which is the case for typical applications. Applying Equation 3 as a layer was previously explored by von Oswald et al. (2023) (who called it a mesa-layer) in the context of understanding deep transformers and by Garnelo and Czarnecki (2023) (who called an intention layer) on the expressivity of self-attention, but here we derive it as mathematical consequence of requiring the sequence model to be able to perform associative recall.
Relationship to linear attention.
Although ${m}_{\textsc{linear},\,t}$ is the optimal linear memory with respect to our objective function, it can be expensive to compute, requiring us to invert a ${D_{k}}\times{D_{k}}$ matrix at every timestep. One crude but hardware-efficient simplification would be to assume that $\mathbf{K}{}^{\top}\mathbf{K}=\mathbf{I}$, avoiding matrix inversion completely, at the cost of some recall ability. After making this rough approximation, we arrive exactly at the equations for linear attention: $\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}\approx\mathbf{V}{}^{\top}\mathbf{K}\mathbf{q}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\mathbf{q}_{t}$.
Our derivation clarifies exactly how linear attention fails as an associative memory. When the keys are zero mean, $\mathbf{K}{}^{\top}\mathbf{K}$ is proportional to the empirical covariance matrix of the keys. Thus linear attention is a crude associative memory that ignores the covariance between the dimensions of the key vectors. In fact, it is an optimal linear memory only when $t\leq{D_{k}}$ and the keys are orthonormal $\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}=\delta_{ij}$.
From a training stability perspective, the inverse covariance factor helps prevent the output $\mathbf{y}_{t}$ from becoming unbounded. We can bound the norm of the output of a least squares layer (Equation 3) by $\lVert\mathbf{y}_{t}\rVert_{2}=\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}\rVert_{2}\leq\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)/\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})$ where $\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})$ is the minimum eigenvalue of $\mathbf{K}{}^{\top}\mathbf{K}$. See Appendix A for the derivation. When we ignore the covariance matrix, approximating it with the identity, the denominator disappears and the output $\mathbf{y}_{t}$ can grow arbitrarily large with the sequence length, leading to training instability. This instability indeed shows up in practice for linear attention, and Qin et al. (2022) proposed output normalization to stabilize training. Our derivation shows that output normalization works by approximately restoring the self-normalizing property of Equation 3 that linear attention loses. We will see in Section 4 that ignoring the covariance causes linear attention to perform worse at associative recall.
Efficient computation.
Conveniently, we can compute $\mathbf{M}_{t}$ in Equation 3 for $t=1,\ldots T$ in parallel or sequentially. To compute them in parallel, we solve a batch of $T$ independent linear systems. To compute them sequentially, we note that $(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}=(\mathbf{K}_{t-1}{}^{\top}\mathbf{K}_{t}+\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})^{-1}$, allowing us to autoregressively compute the inverses via the Woodbury matrix identity. The sequential form is known as online linear regression or recursive least squares (RLS) (Haykin, 2014). Thus during training we can use the parallel form, and during testing we can unroll the recurrence into a recurrent neural network. In prior deep learning literature, this kind of unweighted linear regression layer has been called an intention layer (Garnelo and Czarnecki, 2023) and a mesa-layer (von Oswald et al., 2023), though not motivated by associative memory.
Nonlinear associative memory.
So far we have only considered linear associative memory implemented by linear functions ${\mathcal{M}_{\textsc{linear}}}$ which have limited expressiveness. We can derive nonlinear memories by using a feature map $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$, replacing each key $\mathbf{k}_{i}$ with $\phi(\mathbf{k}_{i})$. Then our function class is defined by $\mathcal{M}_{\phi}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{\phi}}}\}$ and we solve the least-squares objective $\operatorname*{\arg\!\min}_{\mathbf{M}}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\phi(\mathbf{k}_{i})\rVert_{2}^{2}/2$. Solving this least squares objective for $\mathbf{M}$ results in an optimal associative memory that is nonlinear with respect to the input keys. Parameterizing associative memory with a nonlinear feature map has a long history, dating back to Poggio (1975) which found polynomial feature maps to be highly effective, consistent with more recent works (Zhang et al., 2023, Arora et al., 2024).
If we make the same crude approximation as linear attention of disregarding the covariance among $\{\phi(\mathbf{k}_{i})\}_{i=1}^{t}$, we arrive at a suboptimal nonlinear memory. Thus, recent works that apply linear attention with a nonlinear feature map, such as Schlag et al. (2021), Kasai et al. (2021), Zhang et al. (2023), Chen et al. (2024), and others, can be understood within our framework as suboptimal nonlinear memory as well.
Vignette 2: Gated linear attention and state-space models are (suboptimal) weighted linear least squares regression
Classic regression theory tells us that unweighted linear regression works best when the data is stationary (Hastie et al., 2009, Haykin, 2014). However, time-series data and text data are typically non-stationary, and recent tokens are more informative of predicting the next token. Thus a natural requirement is to encourage our memory ${m}_{t}$ to capture more of the recent history by downweighting the past. Fortunately, this can be straightforwardly enforced through classic weighted linear least squares regression together with geometrically decaying weights to forget older key-value pairs.
To develop an architecture from weighted linear least squares, we change only the optimization objective relative to regular linear least squares:
  1. 1.
    Choice of weights: assign time-decaying weights to associations,
    ${m}_{\textsc{linear},\,t}=\operatorname*{\arg\!\min}_{{m}\in{\mathcal{M}_{\textsc{linear}}}}\frac{1}{2}\sum_{i=1}^{t}\gamma_{i}^{(t)}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2},$ (4)
    where the weights are geometrically decaying in time: $\gamma_{i}^{(t)}=\prod_{j=i+1}^{t}\gamma_{j}$ and $\gamma_{i}\in[0,1]$ for $i=1,\ldots,t$.
  2. 2.
    Choice of function class: linear functions ${\mathcal{M}_{\textsc{linear}}}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\mathbf{k},\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}\}$.
  3. 3.
    Choice of minimization algorithm: analytical solution equivalent to one step of Newton’s method.
Here the analytical min-norm solution is
$\mathbf{M}_{t}$ =$=\begin{cases}\mathbf{V}{}^{\top}\bm{\Gamma}\mathbf{K}(\mathbf{K}{}^{\top}\bm{\Gamma}\mathbf{K})^{-1},\ \text{overdetermined, e.g. }\,T\geq{D_{k}}\\ \mathbf{V}{}^{\top}(\mathbf{K}\mathbf{K}{}^{\top})^{-1}\mathbf{K},\ \text{underdetermined, e.g. }\ T<{D_{k}}.\end{cases}$ (5)
where we defined a diagonal matrix $\bm{\Gamma}$ (omitting the $t$ subscript following our convention) with entries $\bm{\Gamma}_{ii}=\gamma_{i}^{(t)}$. The underdetermined case has the same solution in both the weighted and unweighted case.
Relationship to linear attention with gating and state-space models.
Once again, computing $\mathbf{M}_{t}$ requires matrix inversion which is not efficient on matrix-multiplication based hardware. Making the same crude approximation as before of $\mathbf{K}{}^{\top}\bm{\Gamma}\mathbf{K}\approx\mathbf{I}$ results in $\mathbf{M}_{t}\approx\mathbf{V}{}^{\top}\bm{\Gamma}\mathbf{K}=\sum_{i=1}^{t}\gamma_{i}^{(t)}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}$. We can further unroll this equation into a recurrence due to the multiplicative structure of our weights:
$\mathbf{M}_{t}\approx\sum_{i=1}^{t}\gamma_{i}^{(t)}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}=\sum_{i=1}^{t}\left(\prod_{j=i+1}^{t}\gamma_{j}\right)\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}=\gamma_{t}\sum_{i=1}^{t-1}\gamma_{i}^{(t-1)}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}+\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}=\gamma_{t}\mathbf{M}_{t-1}+\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (6)
This is exactly the recurrence implemented by gated variants of linear attention, such as GLA (Yang et al., 2024b), RetNet (Sun et al., 2023), Gateloop (Katsch, 2024), RWKV-6 (Peng et al., 2024), mLSTM (Beck et al., 2024), LRU (Orvieto et al., 2023), as well as state-space models like Mamba-2 (Dao and Gu, 2024). Forgetting older tokens significantly improves performance, which was first discovered empirically (Hochreiter and Schmidhuber, 1997, Greff et al., 2017, van der Westhuizen and Lasenby, 2018). Our derivation of gating from weighted least squares provides a mathematically grounded explanation for the benefits of forgetting through the classic theory of online regression on non-stationary data.
Efficient computation.
Like before, we can compute the memory of our architecture (without approximation) sequentially or in parallel as long as we are willing to invest FLOPs in matrix inversion. Parallel computation can be done by solving a batch of $T$ ${D_{k}}\times{D_{k}}$ linear systems. The sequential version can be derived by applying the Woodbury matrix inversion formula while taking into account our geometrically-decaying weights. Geometrically-weighted recursive least squares (RLS) is a classic method from the adaptive filtering literature (Johnstone et al., 1982).
Vignette 3: Fast weight programmers and online learners are first-order methods for solving streaming least-squares
Up until now we have derived linear associative memory architectures through direct analytical solution, equivalent to one step of Newton’s method, a second-order optimization method. However, second-order iterative methods are computationally expensive, requiring matrix inversions at each step. On the other hand, disregarding the matrix inversion as done by linear attention leads to poor performance, since it ignores the covariance structure of the data.
A computationally cheaper alternative is to apply a first-order gradient-descent method, corresponding to the following design choices:
  1. 1.
    Choice of weights: assign equal weight to each association,
    ${m}_{\textsc{linear},\,t}=\operatorname*{\arg\!\min}_{{m}\in{\mathcal{M}_{\textsc{linear}}}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2}.$ (7)
  2. 2.
    Choice of function class: linear functions ${\mathcal{M}_{\textsc{linear}}}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\mathbf{k},\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}\}$.
  3. 3.
    Choice of minimization algorithm: gradient descent, possibly with adaptive step sizes, momentum, or weight decay (L2 regularization).
We can now derive the test-time regression layer corresponding to a first-order iterative method.
Since we consider linear functions, solving Equation 7 is equivalent to minimizing the objective
$\mathcal{L}_{t}(\mathbf{M})=\sum_{i=1}^{t}L_{i}(\mathbf{M})=\sum_{i=1}^{t}\frac{1}{2}\lVert\mathbf{v}_{i}-\mathbf{M}\mathbf{k}_{i}\rVert_{2}^{2},$ (8)
using its gradient
$\nabla\mathcal{L}_{t}(\mathbf{M})$ $=\sum_{i=1}^{t}\nabla L_{i}(\mathbf{M})=\sum_{i=1}^{t}(\mathbf{M}\mathbf{k}_{i}-\mathbf{v}_{i})\mathbf{k}_{i}{}^{\top}=(\mathbf{M}\mathbf{K}{}^{\top}-\mathbf{V}{}^{\top})\mathbf{K}.$ (9)
to find $\mathbf{M}_{t}$. Let the initialization be at the origin $\mathbf{M}^{(0)}_{t}=\mathbf{0}$. Iteratively solving this objective with full-batch gradient descent with step size $\beta_{i}$ at each iteration results in
$\mathbf{M}^{(i)}_{t}$ $=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\nabla\mathcal{L}_{t}(\mathbf{M}^{(i-1)}_{t})$ (10)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\sum_{j=1}^{t}\nabla L_{j}(\mathbf{M}^{(i-1)}_{t})$ (11)
$=\mathbf{M}^{(i-1)}_{t}-\beta_{i}\sum_{j=1}^{t}(\mathbf{M}^{(i-1)}_{t}\mathbf{k}_{j}-\mathbf{v}_{j})\mathbf{k}_{j}{}^{\top}$ (12)
$=\mathbf{M}^{(i-1)}_{t}(\mathbf{I}-\sum_{j=1}^{t}\beta_{i}\mathbf{k}_{j}\mathbf{k}_{j}{}^{\top})+\beta_{i}\sum_{j=1}^{t}\mathbf{v}_{j}\mathbf{k}_{j}{}^{\top}.$ (13)
After $I$ iterations, we produce the linear memory represented by $\mathbf{M}_{t}=\mathbf{M}^{(I)}_{t}$.
Preconditioning gradient descent.
Consider the $I=1$ case with step size $\beta_{i}=1$. Since we initialize at the origin, we obtain the equation for linear attention
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})=\mathbf{V}{}^{\top}\mathbf{K},$ (14)
a result also shown by (Sun et al., 2024, Irie et al., 2022).
A standard technique to accelerate the convergence of gradient descent is to precondition the gradient (Boyd and Vandenberghe, 2004). The optimal preconditioner in this case, due to the quadratic objective, is the Jacobian of $\mathcal{L}_{t}$, $\mathbf{J}_{\mathcal{L}_{t}}(\mathbf{M})=\mathbf{K}{}^{\top}\mathbf{K}$. This preconditioner allows gradient descent to converge in a single update
$\mathbf{M}^{(1)}_{t}=\mathbf{M}^{(0)}_{t}-\nabla\mathcal{L}_{t}(\mathbf{M}^{(0)}_{t})\mathbf{J}_{\mathcal{L}_{t}}(\mathbf{M}^{(0)}_{t})^{-1}=\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1},$ (15)
reproducing the analytic solution from Equation 3. The preconditioned recurrence is equivalent to one step of Newton’s method, a second order iterative method.
By taking this iterative optimization perspective, we see that linear attention and Equation 3 lie on two extremes: batch gradient descent with no preconditioning versus batch gradient descent with perfect preconditioning. The difference in performance comes from whether the optimizer accounts for the curvature of the objective, governed by the covariance of the keys (Boyd and Vandenberghe, 2004). Exploring other preconditioners, e.g. quasi-Newton methods, that interpolates between the two extremes may be an interesting direction for future research.
Online learners as stochastic gradient descent.
An alternative to full-batch gradient descent is stochastic gradient descent (SGD) where we use a subset of the samples to compute the gradient at each step. This is particularly convenient in causal sequence modeling where keys and values typically are presented one at a time, rather than all at once. In this streaming setting, we can instead perform single-example SGD to iteratively minimize Equation 1. In the case of linear memories ${\mathcal{M}_{\textsc{linear}}}$, Equation 11 with a single-example becomes
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}-\beta_{t}\nabla L_{t}(\mathbf{M}_{t-1})$ (16)
$=\mathbf{M}_{t-1}+\beta_{t}(\mathbf{v}_{t}-\mathbf{M}_{t-1}\mathbf{k}_{t})\mathbf{k}_{t}{}^{\top}$ (17)
$=\mathbf{M}_{t-1}(\mathbf{I}-\beta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})+\beta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top},$ (18)
where $\mathbf{M}_{t}$ is the matrix representation of ${m}_{t}$. This is exactly the recurrence of DeltaNet (Schlag et al., 2021, Yang et al., 2024c). As we perform stochastic gradient descent, we produce an online sequence of memories $\{\mathbf{M}_{i}\}_{i=1}^{t}$, each of which is an associative memory ${m}_{i}$ for key-value pairs up to $(\mathbf{k}_{i},\mathbf{v}_{i})$. The stochastic gradient approach to streaming inputs has the advantage that ${m}_{t}$ focuses more on recent key-value pairs, rather than uniformly over all timesteps as in the case of unweighted RLS and unweighted linear attention. This recency bias allows online learners to handle non-stationary data such as text and time-series.
Equation 18 is known by many names, such as the delta rule or the Widrow-Hoff algorithm (Widrow and Hoff, 1988). Within the adaptive filtering literature this algorithm is known as Least-Mean Squares (LMS) (Haykin, 2014); there it is one of the most common algorithms for handling non-stationary signals. Classic neurocomputing literature (Kohonen, 1989, Equation 6.33) has also previously studied using gradient descent with a single example to update an associative memory map. The iterates $\mathbf{M}_{t}$ are also known as “fast weights”, which were originally proposed by Schmidhuber (1992) as an internal learning/optimization procedure executed at test time, in contrast to the “slow weights” of a neural network that are updated only at training time. It is also possible to apply Equation 16 to iteratively produce nonlinear regressors (Sun et al., 2024).
Adaptive step sizes.
Adaptive step sizes is an intuitive improvement to our gradient-descent-based recurrent architecture. A common way to derive adaptive step sizes in general is by leveraging the relationship between gradient descent and online learning (Cesa-Bianchi et al., 2004, Zinkevich, 2003), a connection that has produced many successful algorithms in the past, such as Adagrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2015). Recently, Liu et al. (2024) derived an effective sequence layer called Longhorn from an online learning perspective (Kulis and Bartlett, 2010), solving
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\ \frac{1}{2}\lVert\mathbf{M}-\mathbf{M}_{t-1}\rVert_{F}^{2}+\frac{\delta_{t}}{2}\lVert\mathbf{v}_{t}-\mathbf{M}\mathbf{k}_{t}\rVert_{2}^{2},$ (19)
resulting in the recurrence
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}\left(\mathbf{I}-\frac{1}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}{\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}}\right)+\frac{1}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (20)
Comparing the Longhorn update in Equation 20 to the SGD update in Equation 18, we see that Longhorn’s layer is equivalent to SGD with an adaptive step size of $\beta_{t}=1/(1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2})$. This alternative perspective reflects the fact that online learning is essentially equivalent to stochastic optimization (Cesa-Bianchi et al., 2004, Duchi et al., 2011, Hazan, 2022).
In Equation 19, $\delta_{t}$ controls how well $\mathbf{M}_{t}$ memorizes the latest association $(\mathbf{k}_{t},\mathbf{v}_{t})$. When $\delta_{t}\to\infty$, Equation 19 is equivalent to a constrained optimization problem
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\ \frac{1}{2}\lVert\mathbf{M}-\mathbf{M}_{t-1}\rVert_{F}^{2}\quad\text{subject to}\ \mathbf{M}_{t}\mathbf{k}_{t}=\mathbf{v}_{t}.$ (21)
Solving this constrained optimization problem results in a step size $\beta_{t}=1/\lVert\mathbf{k}_{t}\rVert_{2}^{2}$ with update
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}\left(\mathbf{I}-\frac{1}{\lVert\mathbf{k}_{t}\rVert_{2}^{2}}{\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}}\right)+\frac{1}{\lVert\mathbf{k}_{t}\rVert_{2}^{2}}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (22)
This special case is also known as the normalized LMS (NLMS) update in adaptive filtering (Castoldi and de Campos, 2009) and known as the projection method (Tanabe, 1971) or Kaczmarz’s method (Kaczmarz, 1937) in optimization theory.
Although here we discuss test-time regressors with adaptive step sizes that only depend on the current timestep, more sophisticated step sizes with long term dependencies, such as Adam (Kingma and Ba, 2015), has been explored by Clark et al. (2022) but are more challenging to parallelize. Given that adaptive optimizers are now a staple of modern deep learning, having come a long way from naive gradient descent, an interesting future line of research would be to integrate better adaptive step sizes into test-time regressors.
L2 regularization and momentum.
In addition to adaptive step sizes, we can add regularization and momentum. We focus on L2 regularization/weight decay here, but concurrent work by Behrouz et al. (2024) derives the case with momentum. Consider L2 regularizing Equation 8 to produce the objective function
$\mathcal{L}_{\text{reg},t}(\mathbf{M})=\sum_{i=1}^{t}L_{i}(\mathbf{M})+\frac{\lambda_{t}}{2}\lVert\mathbf{M}\rVert_{2}^{2}$ (23)
where $\lambda_{t}\in[0,1]$ is the regularization strength. When we minimize this regularized objective with single-example stochastic gradient descent, we obtain the recurrence
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}-\beta_{t}\left[\nabla L_{t}(\mathbf{M}_{t-1})+\lambda_{t}\mathbf{M}_{t-1}\right]$ (24)
$=(1-\beta_{t}\lambda_{t})\mathbf{M}_{t-1}-\beta_{t}\nabla L_{t}(\mathbf{M}_{t-1})$ (25)
$=(1-\beta_{t}\lambda_{t})\mathbf{M}_{t-1}+\beta_{t}(\mathbf{v}_{t}-\mathbf{M}_{t-1}\mathbf{k}_{t})\mathbf{k}_{t}{}^{\top},$ (26)
which is known as Leaky LMS in adaptive filtering (Haykin, 2014). Notably, the first term of our new recurrence resembles Equation 17 but also includes a multiplicative factor similar to a forget gate.
Recently, Yang et al. (2024a) proposed Gated DeltaNet motivated by combining DeltaNet (Schlag et al., 2021) with the forget gating of Mamba-2 (Dao and Gu, 2024), though with a different recurrence $\mathbf{M}_{t}=\mathbf{M}_{t-1}\alpha_{t}\left(\mathbf{I}-\eta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}\right)+\eta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}$. Here we show that the Gated DeltaNet recurrence is actually an equivalent reparameterization of our recurrence in Equation 26:
$\mathbf{M}_{t}$ $=(1-\beta_{t}\lambda_{t})\mathbf{M}_{t-1}+\beta_{t}(\mathbf{v}_{t}-\mathbf{M}_{t-1}\mathbf{k}_{t})\mathbf{k}_{t}{}^{\top}$ (27)
$=\mathbf{M}_{t-1}\left[\underbrace{(1-\beta_{t}\lambda_{t})}_{\alpha_{t}}\mathbf{I}-\beta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}\right]+\beta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}$ (28)
$=\mathbf{M}_{t-1}\alpha_{t}\left[\mathbf{I}-\underbrace{\frac{\beta_{t}}{\alpha_{t}}}_{\eta_{t}}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}\right]+\frac{\beta_{t}}{\alpha_{t}}\underbrace{(\alpha_{t}\mathbf{v}_{t})}_{\mathbf{v}^{\prime}_{t}}\mathbf{k}_{t}{}^{\top}$ (29)
[…]
Thus Gated DeltaNet is equivalent to coupling L2 regularization with the step size and rescaling $\mathbf{v}_{t}$ by $\alpha_{t}$. We can transform Equation 30 into Equation 26 via the invertible mapping $(\alpha_{t},\eta_{t},\mathbf{v}_{t}^{\prime})=(1-\beta_{t}\lambda_{t},\beta_{t}/(1-\beta_{t}\lambda_{t}),(1-\beta_{t}\lambda_{t})\mathbf{v}_{t})$. The inverse map is given by $(\lambda_{t},\beta_{t},\mathbf{v}_{t})=((1-\alpha_{t})/(\alpha_{t}\eta_{t}),\alpha_{t}\eta_{t},\allowbreak\mathbf{v}^{\prime}_{t}/\alpha_{t})$. Past works have shown that decoupling weight decay from the step size results in better generalization and easier-to-tune hyperparameters (Loshchilov and Hutter, 2019), suggesting that similar modifications to Gated DeltaNet may improve model performance.
Vignette 4: Softmax attention is nonparametric regression
Having exhaustively derived multiple associative memory architectures via linear and nonlinear parametric regression, we can now derive self-attention-like architectures using the tools of nonparametric regression.
Unnormalized softmax attention is (suboptimal) kernel regression.
We start first from kernel regression as most common nonparametric regression method (Schölkopf and Smola, 2001), applying the kernel trick to our feature-mapped linear memory. Let $k:\mathbb{R}^{{D_{k}}\times{D_{k}}}\to\mathbb{R}$ be a Mercer kernel such that $k(\mathbf{k}_{i},\mathbf{k}_{j})=(\Phi\Phi{}^{\top})_{ij}$ where $\Phi=\begin{bmatrix}\phi(\mathbf{k}_{1})\ldots\phi(\mathbf{k}_{t})\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{\phi}}}$ and $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$ is an infinite-dimensional feature map. Following our recipe for model design, we make the following design choices:
  1. 1.
    Choice of weights: assign equal weight to each association
    ${m}_{k,\,t}=\operatorname*{\arg\!\min}_{{m}\in\mathcal{M}_{k}}\frac{1}{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}-\mathbf{M}\phi(\mathbf{k}_{i})\rVert_{2}^{2},$ (31)
    where $\phi$ is an infinite-dimensional feature map.
  2. 2.
    Choice of function class: linear functions $\mathcal{M}_{k}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}\}$.
  3. 3.
    Choice of minimization algorithm: analytical solution.
The output of the optimal kernelized nonlinear associative memory induced by kernel $k$ at timestep $t$ is:
${m}_{k,\,t}(\mathbf{q})=\mathbf{V}{}^{\top}k(\mathbf{K},\mathbf{K})^{-1}k(\mathbf{K},\mathbf{q}),$ (32)
equivalent to kernel regression. Here we follow standard shorthands from the kernel learning literature: $k(\mathbf{K},\mathbf{K})$ is a shorthand for the matrix with $k(\mathbf{k}_{i},\mathbf{k}_{j})$ in entry $(i,j)$; $k(\mathbf{K},\mathbf{q})$ is a shorthand for the vector with $k(\mathbf{k}_{i},\mathbf{q})$ in entry $i$. Interestingly, when we use an exponential kernel $k(\mathbf{k}_{i},\mathbf{k}_{j})=\exp(\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}/\sqrt{{D_{k}}})$, corresponding to an infinite-dimensional feature map, and make the crude approximation $k(\mathbf{K},\mathbf{K})=\mathbf{I}$, we obtain an associative memory equivalent to an unnormalized softmax attention. However unnormalized softmax attention is empirically known to be unstable to train (Lu et al., 2021). Thus we do not stop our derivations here; instead, we will see how the standard normalized softmax attention comes from another kind of nonparametric regressor.
Softmax attention with QKNorm is a locally-constant nonparametric regressor.
Local polynomial regressors is another classic class of nonparametric estimators (Wasserman, 2006), which have been historically used as nonlinear smoothing functions in data analysis (Fan, 2018). Given a query $\mathbf{q}\in\mathbb{R}^{D_{k}}$ at time $t$ that we would like to use to retrieve memory/make a prediction, let our function class be $\mathcal{M}_{\mathbf{q}}^{p}$, the set of all vector-valued functions that are locally a degree $p$ polynomial around points similar to $\mathbf{q}$, as measured by some similarity function $s$ that only depends on the distance to $\mathbf{q}$, i.e. $s(\mathbf{k},\mathbf{q})=s(\mathbf{k}-\mathbf{q})$. A classic example is the exponential smoothing kernel $s_{\text{exp},\,B}(\mathbf{k},\mathbf{q})=\exp(-\lVert\mathbf{k}-\mathbf{q}\rVert_{2}^{2}/B)$ where $B$ is a bandwidth parameter controlling the local neighborhood size around $\mathbf{q}$. Typically local polynomial regression is done for 1-dimensional data; here we generalize it to the multivariate case.
Nonparametric regression with local polynomial estimators corresponds to the following design choices:
  1. 1.
    Choice of weights: assign weights to associations based on similarity to $\mathbf{q}$,
    ${m}_{(p,\mathbf{q}),\,t}=\operatorname*{\arg\!\min}_{{m}\in\mathcal{M}^{p}_{\mathbf{q}}}\frac{1}{2}\sum_{i=1}^{t}\gamma_{i}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2}$ (33)
    with weights given by $\gamma_{i}=s(\mathbf{k}_{i},\mathbf{q})=s(\mathbf{k}_{i}-\mathbf{q})$. We choose weights that are monotonic in distance to ensure that ${m}_{(p,\mathbf{q}),\,t}$ is an accurate approximation within a local neighborhood of $\mathbf{q}$ defined by $s$.
  2. 2.
    Choice of function class: $\mathcal{M}_{\mathbf{q}}^{p}$, the class of polynomials around $\mathbf{q}$, such that any ${m}\in\mathcal{M}_{\mathbf{q}}^{p}$ can be written as:
    ${m}(\mathbf{k}_{i})=\mathbf{M}^{(0)}+\mathbf{M}^{(1)}(\mathbf{k}_{i}-\mathbf{q})+\mathbf{M}^{(2)}(\mathbf{k}_{i}-\mathbf{q},\mathbf{k}_{i}-\mathbf{q})+\ldots+\mathbf{M}^{(p)}(\underbrace{\mathbf{k}_{i}-\mathbf{q},\ldots,\mathbf{k}_{i}-\mathbf{q}}_{p\ \text{arguments}})$ (34)
    where $\mathbf{M}^{(j)}$ is an order $j$ multilinear map [^4] See e.g. lecture notes by Conrad for the multivariate Taylor expansion case.. This can be thought of as an order $p$ Taylor expansion approximation of the true key-value mapping around $\mathbf{q}$.
  3. 3.
    Choice of minimization algorithm: analytical solution. Solving for the minimum is equivalent to solving for the set of tensors $(\mathbf{M}^{(0)},\ldots,\mathbf{M}^{(p)})$ that minimize the least squares objective. Our prediction on $\mathbf{q}$ is then ${m}_{(p,\mathbf{q}),\,t}(\mathbf{q})=\mathbf{M}^{(0)}$.
Unfortunately, solving Equation 33 is computationally expensive for large $p$. Instead, we can restrict ourselves to the simplest case of $p=0$, corresponding to the class of locally constant functions around $\mathbf{q}$, also known as Nadaraya-Watson estimators. With this simplification, the output of our nonparametric associative memory is
$m_{(0,\mathbf{q}),\,t}(\mathbf{q})$ $=\operatorname*{\arg\!\min}_{\mathbf{M}^{(0)}\in\mathbb{R}^{D_{k}}}\frac{1}{2}\sum_{i=1}^{t}s(\mathbf{k}_{i},\mathbf{q})\lVert\mathbf{v}_{i}-\mathbf{M}^{(0)}\rVert_{2}^{2}=\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i}.$ (35)
We may be disappointed at first that $s$ should be a function of the distance $\lVert\mathbf{k}-\mathbf{q}\rVert_{2}$ to enforce the local approximation property of the estimator in Equation 33; in contrast, softmax attention relies on an exponentiated inner product. However, if we normalize keys and queries such that $\lVert\mathbf{k}\rVert_{2}=\lVert\mathbf{q}\rVert_{2}=1$, the exponential smoothing kernel with bandwidth $B=2\sqrt{{D_{k}}}$ is equivalent to the exponential scaled dot product:
$s_{\text{exp},2\sqrt{D_{k}}}(\mathbf{k},\mathbf{q})=\exp\left(\frac{-\lVert\mathbf{k}-\mathbf{q}\rVert_{2}^{2}}{2\sqrt{{D_{k}}}}\right)=\exp\left(\frac{-\lVert\mathbf{k}\rVert_{2}^{2}-\lVert\mathbf{q}\rVert_{2}^{2}+2\mathbf{k}{}^{\top}\mathbf{q}}{2\sqrt{{D_{k}}}}\right)\propto\exp\left(\frac{\mathbf{k}{}^{\top}\mathbf{q}}{\sqrt{{D_{k}}}}\right),$ (36)
allowing us to exactly derive softmax attention in Equation 35. Moreover, this preprocessing step lines up with an existing technique called query-key normalization (QKNorm), empirically known to stablize the training of large transformers (Dehghani et al., 2023, Wortsman et al., 2023) but lacks theoretical justification. Our derivation provides a simple explanation: softmax attention with QKNorm is a locally zeroth order non-parametric approximation of the true key-value map.
Figure 2: Solving a non-stationary online regression task without any learnable parameters. The regression-memory correspondence enables us to perform regression via a single forward pass of a test-time regression layer. The inputs change more rapidly in initial $T/4$ timesteps, making it harder for regressors to learn. The latter $43T/$ 4 timesteps are more stable; however, without decaying weights, linear attention and unweighted RLS are unable to adjust to the resulting non-stationary data.
Higher-order attention.
Equation 33 suggests a path to higher-order generalizations of self-attention by considering $p>0$ at the cost of more computation. By reducing Equation 33 to a weighted least squares problem, we see the $p=1$ case of higher-order softmax self-attention is
$\begin{bmatrix}\mathbf{M}^{(0)}&\mathbf{M}^{(1)}\end{bmatrix}$ $=(\mathbf{V}{}^{\top}\mathbf{W}\mathbf{X})(\mathbf{X}{}^{\top}\mathbf{W}\mathbf{X})^{-1},$ $\mathbf{W}$ $\triangleq\operatorname{diag}\left(s(\mathbf{k}_{1},\mathbf{q}),\ldots,s(\mathbf{k}_{t},\mathbf{q})\right)$ (37)
$\mathbf{X}$ $=\begin{bmatrix}1&\ldots&1\\ \mathbf{k}_{1}-\mathbf{q}&\ldots&\mathbf{k}_{t}-\mathbf{q}\end{bmatrix}{}^{\top}$ $\mathbf{V}$ $=\begin{bmatrix}\mathbf{v}_{1}&\ldots&\mathbf{v}_{t}\end{bmatrix}{}^{\top},$ (38)
where we recall that $\mathbf{M}^{(0)}\in\mathbb{R}^{{D_{v}}}$ and $\mathbf{M}^{(1)}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}$. An alternate form for the prediction $\mathbf{M}^{(0)}$ that is more intuitive is
$\mathbf{M}^{(0)}$ $=\bar{\mathbf{v}}-\mathbf{M}^{(1)}\bar{\mathbf{d}},$ $\bar{\mathbf{v}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i},$ $\bar{\mathbf{d}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{d}_{i},$ $\mathbf{d}_{i}$ $\triangleq\mathbf{k}_{i}-\mathbf{q},$ (39)
resulting in an offset softmax attention as the output: $m_{(1,\mathbf{q}),\,t}(\mathbf{q})=\mathbf{M}^{(0)}=\bar{\mathbf{v}}-\mathbf{M}^{(1)}\bar{\mathbf{d}}$. This alternate form comes from the fact that if we already know $\mathbf{M}^{(1)}$, then we can solve the $p=0$ case with an alternate regression problem with preprocessed data $\{(\mathbf{k}_{i},\mathbf{v}_{i}-\mathbf{M}^{(1)}\mathbf{d}_{i})\}_{i=1}^{t}$, resulting in $\mathbf{M}^{(0)}$ being a normalized weighted sum of $\mathbf{v}_{i}-\mathbf{M}^{(1)}\mathbf{d}_{i}$. Unlike standard softmax attention (the $p=0$ case), which only considers the interaction between keys and queries, the $p=1$ case in Equation 37 also accounts for the covariance between keys through the inverse term, similar to how linear regression improves upon linear attention from before.
Unfortunately, directly computing Equation 37 for $\mathbf{q}\in\{\mathbf{q}_{1},\ldots,\mathbf{q}_{T}\}$ in parallel to fully leverage hardware-acceleration is impractical for all but the smallest sequence problems. Materializing $\{\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ space, and materializing $\{\mathbf{X}_{t}{}^{\top}\mathbf{W}_{t}\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T{D_{k}}^{2}\right)$ space. Computing $\{\mathbf{X}{}^{\top}_{t}\mathbf{W}_{t}\allowbreak\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ time, and solving a batch of $T$ linear systems requires $\mathcal{O}\left(T{D_{k}}^{3}\right)$ time. These challenges makes the current form computationally impractical for all but the smallest sequence problems. Developing hardware-aware optimizations to compute Equation 37 similar to FlashAttention (Dao et al., 2022) is a promising future direction that would make higher order attention practical.
We compare our $p=1$ generalization of softmax attention against three other models on a non-stationary online regression problem. The keys are generated via a non-stationary (switching) autoregressive process:
$\mathbf{k}_{t+1}=\begin{cases}0.9\mathbf{k}_{t}+0.02\bm{\varepsilon}_{t},\,t<T/4\\ 0.999\mathbf{k}_{t}+0.02\bm{\varepsilon}_{t},\,t\geq T/4\end{cases}$ (40)
with ${D_{k}}=64$ and $T=256$. Each $\bm{\varepsilon}_{t}\sim\operatorname{Normal}\left(0,\bm{\Sigma}\right)$ where $\bm{\Sigma}$ has close to unit diagonals and non-zero off-diagonals. The choice of a non-identity $\Sigma$ creates instantaneous correlation between the dimensions of the keys so that $\mathbf{K}{}^{\top}\mathbf{K}\neq\mathbf{I}$. We let $\mathbf{v}_{t}=\lVert\mathbf{k}_{t+1}\rVert_{2}$, resulting in a next-token prediction task. Through the regression-memory correspondence, we directly apply each of these test-time regression layers to the key-value pairs without any learnable parameters; we simply pass the sequences through each layer via a single forward pass. We plot the one-step out-of-sample test loss $\lVert\mathbf{v}_{t+1}-{m}_{t}(\mathbf{k}_{t+1})\rVert_{2}^{2}$ in Figure 2. During the initial $T/4$ timesteps, the dynamical system has a faster decay time with more rapid changes, making it less predictable. In the latter $43T/$ 4 timesteps, the system slows down and is more predictable. However, unweighted recursive least squares is unable to adapt to this change since it cannot discount datapoints from before the transition; as a result, it maintains a high error throughout the latter timesteps. In contrast, the nonparametric regressors are able to adapt quickly to the transition since earlier keys are dissimilar to the later keys. Interestingly, increasing our nonparametric softmax regressor from $p=0$ to $p=1$ allows the model to better adapt to both the initial fast-changing period and the sudden regime change. As expected, linear attention performs the worst since $\mathbf{K}{}^{\top}\mathbf{K}\neq\mathbf{I}$ in our dynamical system [^5] We used a standard output normalization for linear attention (Qin et al., 2022) to prevent the output from scaling linearly with the sequence length..
Section 4 Constructing effective key-value pairs
Having shown that models derived via regression are able to perform associative recall, we now discuss the importance of constructing test-time key-value pairs that are pertinent to the task at hand. This echoes an unchanging principle in machine learning: a well-designed model is only as effective as the data it processes. Historically, query-key-value sequences were constructed via a linear projection on the corresponding timestep: $\mathbf{q}_{t}=\mathbf{W}_{Q}\mathbf{x}_{t},\,\mathbf{k}_{t}=\mathbf{W}_{K}\mathbf{x}_{t},\,\mathbf{v}_{t}=\mathbf{W}_{V}\mathbf{x}_{t}$ (Vaswani et al., 2017). Our framework naturally generalizes to the multihead case (Shazeer, 2019, Ainslie et al., 2023), corresponding to setting up parallel test-time regression problems, possibly with shared inputs or outputs.
Recent works on recurrent models, starting with Fu et al. (2022), have identified the importance of applying a “short convolution” to construct the query-key-value sequences, which can express the “shift” operation that induction heads learn (Olsson et al., 2022, von Oswald et al., 2023, Fu et al., 2022). The short convolution is a crucial component of purely recurrent language models, resulting in severe performance drops if removed (Yang et al., 2024b, c, a, Sun et al., 2024). In the case of transformers or model that combine recurrence with softmax attention (De et al., 2024), an explicit short convolution is not needed since earlier softmax attention layers can learn the behavior of a short convolution as empirically observed by von Oswald et al. (2023). Here we show that this short convolution is crucial to associative recall, creating bigram-like key value pairs.
One short conv is all you need for associative recall.
Consider an abstracted version of multi-query associative recall (MQAR), a standard associative recall task introduced by Arora et al. (2023), in which a model must be able to recall multiple key-value pairings. In this task, there are $P$ pairs of unique and consistent cue-response [^6] We use the “cue” and “response” terminology to avoid confusion with “keys” and “values” in the context of test-time regression. tokens $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}_{j\in[P]}$ with each cue $\mathbf{u}_{j}$ mapping one-to-one to its response $\mathbf{v}_{j}$. The model is given a contextual sequence of randomly drawn, possibly repeated, cue-response pairs followed by a previously seen cue: $(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T-1},\mathbf{x}_{T},\mathbf{x}_{T+1})=(\mathbf{u}_{i_{1}},\mathbf{v}_{i_{1}},\ldots,\mathbf{u}_{i_{T/2}},\mathbf{v}_{i_{T/2}},\mathbf{u}_{i_{j}})$. $\mathbf{u}_{i_{j}}$ is a cue that appeared earlier (the $j$th pair) and the model is expected to output the paired response $\mathbf{v}_{i_{j}}$ at timestep $T+1$.
We now show that a single test-time regression layer (e.g. as simple as linear attention) with one short convolution is sufficient to solve MQAR, without any parameters other than the embeddings, as long as it is provided with the appropriate key-value pairs to regress. To solve this task, it suffices to test-time memorize all the bigram pairing $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}$. Given $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$, we causally construct our test-time regression dataset as follows. Let the keys be constructed via a short convolution $\mathbf{k}_{t}=w_{0}\mathbf{x}_{t}+w_{1}\mathbf{x}_{t-1}=\mathbf{x}_{t-1}$ where $w_{0}=0,w_{1}=1$. Let the queries and values be the same as the inputs: $\mathbf{q}_{t}=\mathbf{v}_{t}=\mathbf{x}_{t}$. At timestep $T+1$, we have $\mathbf{q}_{T+1}=\mathbf{x}_{T+1}=\mathbf{u}_{i_{j}}$. The output of a linear attention layer is then $\mathbf{y}_{T+1}=\sum_{t=1}^{T+1}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}\mathbf{q}_{T+1}=\sum_{t=1}^{T+1}\mathbf{x}_{t}\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}$ When the embedding space is large enough, we can set the embeddings such that all tokens are orthonormal. Then $\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}$ is nonzero only when $\mathbf{x}_{t-1}=\mathbf{u}_{i_{j}}$, simplifying the output to a sum of only tokens that are followed by $\mathbf{u}_{i_{j}}$. Since the cue-response map is one-to-one, the remaining tokens are all $\mathbf{v}_{i_{j}}$, producing the output $\mathbf{y}_{T+1}\propto\mathbf{v}_{i_{j}}$, solving the task.
Memory capacity, not sequence length, limits model performance.
Figure 3: A single test-time regression layer with one short convolution suffices to solve MQAR with $P=64$ cue-response pairs. When a model has sufficient capacity to memorize all $P$ pairs, it can solve MQAR regardless of the sequence length. The regression problem is stationary (the cue-response map doesn’t change over time), hence no forgetting mechanism is needed to solve the task.
Typical evaluations of models on the MQAR task look at model performance with respect to the length of the sequence, possibly varying the model capacity. A few examples include Figure 2 of Arora et al. (2023), Figure 8 of Dao and Gu (2024), Figure 2 of Yang et al. (2024c), and likely others. However, our above construction shows that the sequence length doesn’t matter for memorizers with a large enough memory capacity; once the memory is large enough, the difficulty of MQAR is independent of the sequence length $T$, which may be counter to the intuitive notion of longer sequences being more difficult. Instead, models are limited by their ability to fully memorize all $P$ key-value pairs.
Although our construction relied on the embedding space being large enough to allow orthonormality, trained models can still succeed without perfectly orthonormal embeddings, which we empirically show in Figure 3. We train on the MQAR task following the procedure of (Arora et al., 2023). However, we use only one test-time regression layer and only use a single size 2 convolution for constructing the keys, following our construction. We also remove the typical MLP block following the sequence layer, since our construction shows that additional nonlinearities are unnecessary for this synthetic associative recall task. We fix our key dimensions ${D_{k}}$ to be the same as the embedding size, $d_{\text{model}}=64$. We compare two test-time regression architectures, linear attention and unweighted recursive least squares (RLS) of Equation 3. We do not use gated variants of linear attention since this task does not require any forgetting; the cue-response mapping is unchanged throughout the sequence. As shown by our previous construction, even linear attention is able to solve the task perfectly when the appropriate keys are constructed. In the case where $d_{\text{model}}=P=64$, linear attention solves MQAR perfectly, independent of sequence length as predicted. Once the number of keys increases to $P=128$, linear attention can no longer perfectly solve MQAR, since it is impossible for 128 keys to be orthogonal in a $d_{\text{model}}=64$ dimensional space. In contrast, an unweighted RLS layer improves upon linear attention by accounting for the key covariance matrix $\mathbf{K}{}^{\top}\mathbf{K}$, allowing it to better test-time memorize the cue-response map, even when $P>d_{\text{model}}$. As a result, RLS is able to solve MQAR regardless of the sequence length.
Section 5 Conclusion and discussion
We have presented a unifying framework that views sequence models with associative memory as test-time regressors. Each architecture emerges naturally by specifying three aspects: the relative importance of each association, the function class, and the optimization algorithm. Through this lens, we derived linear attention variants as approximations to recursive least squares, online learners as first-order least-squares solvers, and softmax attention as a local polynomial regressor. This perspective provided principled explanations for architectural choices like QKNorm and derived higher-order generalizations of attention that go beyond pairwise interactions. We also demonstrated that for multi-query associative recall (MQAR), a single short convolution and a sequence layer suffices, in contrast to typical approaches that use multiple layers and unnecessary feedforward blocks (Arora et al., 2023).
Our paper focuses specifically on sequence architectures that use query-key-value patterns for associative recall. Several important architectural families lie outside this scope, including structured state-space models with more general structured masks (Dao and Gu, 2024), convolutional architectures (Shi et al., 2023, Poli et al., 2023, Gehring et al., 2017, Gu et al., 2021, Bradbury et al., 2017, Smith et al., 2023, Hochreiter and Schmidhuber, 1997), and more novel computational patterns (Qin et al., 2023, Ren et al., 2023). Moreover, complementary research on model backbones and initialization continues to yield important advances and performance improvements (Gu and Dao, 2024, Dao and Gu, 2024, Orvieto et al., 2023, Mehta et al., 2023, Hua et al., 2022).
Our framework opens rich directions for future research by connecting to the extensive literature on regression and optimization. While initial successes like TTT-MLP (Sun et al., 2024) hint at the potential of neural test-time regressors, the space of nonlinear neural regression models remains largely unexplored. Recent work integrating weight decay and momentum (Behrouz et al., 2024, Yang et al., 2024a) suggests opportunities for developing more effective, stable, and parallelizable test-time optimizers. Concurrent work by Behrouz et al. (2024) also suggests opportunities to more creatively utilize associative memory. The practical success of these approaches will depend heavily on efficient hardware implementations (Hua et al., 2022, Yang et al., 2024c, Sun et al., 2024, Dao and Gu, 2024, Dao et al., 2022), aligning with established compute-scaling principles (Sutton, 2019, Hoffmann et al., 2022, Kaplan et al., 2020).
In the words of Kohonen (1989), “associative memory is a very delicate and complex concept which often has been attributed to the higher cognitive processes, especially those taking place in the human brain”. We speculate that test-time associative memory, along with test-time compute more broadly (Sun et al., 2024, 2020, Akyürek et al., 2024), will be fundamental to developing truly adaptive models that can update and learn in changing environments (OpenAI, 2024).
Acknowledgements.
This work was supported in part by AFOSR Grant FA9550-21-1-0397, ONR Grant N00014-22-1-2110, NSF Grant 2205084, the Stanford Institute for Human-Centered Artificial Intelligence (HAI). EBF is a Chan Zuckerberg Biohub – San Francisco Investigator.
Appendix A Bounding the norm of a linear regression layer
Let $\mathbf{y}_{t}=\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}$ and let $\lVert\mathbf{A}\rVert_{2}$ indicate the spectral norm of a matrix $\mathbf{A}$, a matrix norm induced by the L2 norm on vectors. Then
$\lVert\mathbf{y}_{t}\rVert_{2}$ $=\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}\rVert_{2}$ (41)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\sup_{\lVert\mathbf{q}\rVert_{2}=1}\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}\rVert_{2}$ (42)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (43)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}{}^{\top}\mathbf{K}\rVert_{2}\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (44)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\rVert_{2}\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (45)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (46)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\lambda_{i}((\mathbf{K}{}^{\top}\mathbf{K})^{-1})$ (47)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\frac{1}{\lambda_{i}(\mathbf{K}{}^{\top}\mathbf{K})}$ (48)
$=\frac{\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}}{\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}.$ (49)
Hence $\lVert\mathbf{y}_{t}\rVert_{2}\leq\frac{\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}}{\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}$. When we approximate $\mathbf{K}{}^{\top}\mathbf{K}\approx\mathbf{I}$, as in linear attention, we lose the self-normalizing property of dividing by ${\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}$ since the denominator becomes 1. This explains how output normalization by Qin et al. (2022) is an attempt at restoring this intrinsic self-normalizing property of linear regression.